OpenStudy (anonymous):

The function f(t) = 7 cos(pi over 4t) + 12 represents the tide in Light Sea. It has a maximum of 19 feet when time (t) is 0 and a minimum of 5 feet. The sea repeats this cycle every 8 hours. After two hours, how high is the tide? 12 feet 16 feet 8.5 feet 16.5 feet

OpenStudy (anonymous):

\[f(t)=7\cos(\pi/4 t+12)\]

OpenStudy (anonymous):

The equation that you have written down seems incorrect. I think it should be \[f(t) = 7\cos (\pi/4 \times t) +12\] Attached is the graph of the equation. The answer is 12 feet

OpenStudy (anonymous):

Oops! Here's the graph.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Your welcome